Fermi National Laboratory


Accelerator Update

The integrated luminosity for last week was 8965 nb-1.

Monday January 26
The day shift began with Operations monitoring store 3191, with the Antiproton Source (Pbar) stacking, and with MiniBooNE taking beam.

At 8:48 AM, the store aborted due to a LCW pressure problem that caused all the Tevatron (TeV) RF stations to trip off.

A Booster power supply (GMPS) tripped off at 8:50 AM. It tripped off two more times over the next hour.

Operations put the TeV through a dry squeeze at 8:58 AM.

At 9:02 AM, operators made the Collider Detector (CDF) and D-zero (D0) collision halls ready for experimenter access.

The TeV suffered a quench at sectors DA and BA at 9:18 AM.

At 10:26 AM, cryo system technicians began repairs on the TeV sector E2 wet engine.

Both CDF and D0 were out of access by 10:54 AM.

At 11 AM, the cryo techs found a bad expander piston seal on the E2 wet engine. They estimate a five-hour repair to overhaul the engine.

Pbar experts began a reverse proton study and Booster experts began a collimator study around noon. The Booster study halted beam to MiniBooNE.

At 1:35 PM, GMPS tripped off again. Experts replaced a current meter.

A Main Injector (MI) expert began a slipstacking study at 1:37 PM.

D0 came out of access at 2:42 PM and the CDF area, after some interlock repairs, was made up by 3:30 PM. CDF experimenters then went back in on access until 6 PM.

Facility Services personnel installed a generator for the Meson area at MS1 to power the interlock system. This area will be without power until Wednesday, but the generator will only stay until the interlock battery backup gets recharged.

The Booster study finished at 5 PM. Operations resumed sending beam to MiniBooNE. v Pbar studies finished at 5:46 PM.

Operations put the TeV through a dry squeeze at 8:21 PM and then into shot setup at 8:47 PM.

At 11:14 PM, Operations established store 3195 with an initial luminosity of 43.37E30. Pbar resumed stacking.

Operations reported at 11:50 PM that the cryo vacuum at TeV sector E1 was failing. Since the store would abort if the magnets warmed, Cryo techs manually operated the system to try and keep up with the increasing heat load.

Tuesday January 27
The midnight shift began with Operations monitoring store 3195, with cryo techs manually operating the E1 cryo system, with Pbar stacking, and with MiniBooNE taking beam.

At 1:38 AM, the E1 cold compressor tripped off. Operators dumped the store before it had a chance to quench.

Operators prepared the TeV for an access at 2 AM.

Vacuum techs accessed the TeV around 4 AM. At 6 AM, they reported replacing a vacuum pump at sector E13. Cryo techs at CHL began cooling the system.

At 9:49 AM, Operations put the TeV into a forty-minute dry squeeze. At 11:54 AM, Operations put the TeV through a twenty-minute wet squeeze.

Operations put the TeV into shot setup at 12:42 PM and at 2:45 PM established store 3197. Its initial luminosity was 40.86E30. Pbar resumed stacking.

The High Voltage Expert energized feeder 30 at 3:36 PM.

At 5:17 PM, Operations set MI to no ramp for a Recycler study.

The Tevatron Electron Lens (TEL) tripped off on an over temperature indication at 9:57 PM. Operators had it back on five minutes later. It tripped off again at 10:48 PM. Operators found that an internal fan wasn't working and called an expert for advice. Operators stationed a big fan on the power supply to keep it cool until repairs could be made.

A MI RF station (MIRF13) had been tripping off all evening. At 11:12 PM, operators turned off the station and then made adjustments to allow the system to work with seventeen stations.

Wednesday January 28
The midnight shift began with Operations monitoring store 3197, with MiniBooNE taking beam, and with Pbar stacking.

At 12:37 AM, a MI RF station (MIRF10) tripped off the first of eight times over the next two hours.

MI RF stations 10, 15, 17, and 18 all tripped off at 2:33 AM. They all reset okay.

At 2:50 AM, MI experts descended on the Main Control Room (MCR) to conduct coalescing studies.

At 4:05 AM, an operator noticed and resolved a MI clock event conflict.

MIRF10 tripped off at 3 AM and then again at 4:20 AM.

At 4:46 AM, operators bypassed MIRF13 and reconfigured the RF system. Techs will repair this station and MIRF10 later in the morning.

Today's Plans
The plan for today is to allow Linac experts to access their tunnel later this morning. Booster and MI will also go in on access.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 1/28/2004   email Fermilab